OpenStudy (anonymous):

What multiplies to 168 and adds to 74?

OpenStudy (anonymous):

[x ^{2}+74x+168=0

OpenStudy (shamil98):

xy = 168 x + y =74 you could solve it like this, kinda tedious tho

OpenStudy (isaiah.feynman):

@shamil98 You have taught me something new!!!

OpenStudy (shamil98):

lol, glad to be of some help, i'm still thinking of the numbers from guessing tho

OpenStudy (anonymous):

I feel like iI've guessed them all, and because it's online I know that no solution is not the answer

OpenStudy (alekos):

there are no two integers that will do this as far as i can tell

OpenStudy (isaiah.feynman):

I'm in uni, but when factoring I'm lucky to come up with numbers within the multiplication tables.

OpenStudy (shamil98):

yeah the values you get are not whole numbers

OpenStudy (shamil98):

do you have answer choices perhaps?

OpenStudy (alekos):

using the quadratic formula you dont get integers

OpenStudy (anonymous):

the original problem was: \[\log _{4}(x+13)=5-\log _{5}(x+61)\]

OpenStudy (shamil98):

Ah.

OpenStudy (shamil98):

\[\huge \log_4 (x+13) + \log_5(x+61) = 5\] they aren't the same bases so we can't use that log rule...

OpenStudy (anonymous):

that should be a 4 not 5 in front of the (x+61)

OpenStudy (shamil98):

oh. that makes it easier lol

OpenStudy (anonymous):

my bad

OpenStudy (shamil98):

\[\huge \log_4(x+13) + \log_4(x+61) = 5\] using the rule \[\huge \log_bM + \log_bN = \log_b(MN)\]

OpenStudy (shamil98):

\[\huge \log_4 (x+13)(x+61) = 5\]

OpenStudy (shamil98):

And now it's pretty easy.

OpenStudy (anonymous):

right, that's when I got \[x ^{2}+74x+793=625 \]

OpenStudy (shamil98):

Let (x+13)(x+61) = a \[\huge \log_4 a = 5\] using the log rule \[\huge \log_a x = N --> a^n = x\]

OpenStudy (anonymous):

which got me to where I am now

OpenStudy (shamil98):

\[\huge (x+13)(x+61) = 4^5\]

OpenStudy (shamil98):

\[\huge (x+13)(x+61) = 1024\]

OpenStudy (anonymous):

hahaha well im stupid

OpenStudy (shamil98):

I think you calculuated wrong lol

OpenStudy (anonymous):

I most definitely did, thank you so much!!

OpenStudy (shamil98):

Anytime! Glad to be of help.

OpenStudy (isaiah.feynman):

shamil was faster than me!

OpenStudy (shamil98):

lol, I actually learned logs quite recently so it's still fresh in my mind xD

OpenStudy (isaiah.feynman):

I mean typing!

OpenStudy (shamil98):

Oh, that, yeah I got used to writing latex hehe

OpenStudy (isaiah.feynman):

Still in high school?

OpenStudy (shamil98):

Yeah.

OpenStudy (shamil98):

I'm in pre-calc, atm, I skipped algebra 2 so I had to learn logs for my class lol, my calculus adventures are just goals for myself , if you're wondering.

OpenStudy (isaiah.feynman):

Cool. I learned logs long ago...I just utilize their rules in university calculus.